### v2

# thanks: Corresponding author

Abstract
Real-time and accurate prediction of aerodynamic flow fields around airfoils is crucial for flow control and aerodynamic optimization. However, achieving this remains challenging due to the high computational costs and the non-linear nature of flow physics. Traditional Computational Fluid Dynamics (CFD) methods face limitations in balancing computational efficiency and accuracy, hindering their application in real-time scenarios. To address these challenges, this study presents AeroDiT, a novel surrogate model that integrates scalable diffusion models with transformer architectures to address these challenges. Trained on Reynolds-Averaged Navier-Stokes (RANS) simulation data for high Reynolds-number airfoil flows, AeroDiT accurately captures complex flow patterns while enabling real-time predictions. The model demonstrates impressive performance, with average relative $L_{2}$ errors of 0.1, 0.025, and 0.050 for pressure $p$ and velocity components $u_{x},u_{y}$, confirming its reliability. The transformer-based structure allows for real-time predictions within seconds, outperforming traditional U-net diffusion models. This work underscores the potential of generative machine learning techniques to advance computational fluid dynamics, offering potential solutions to persistent challenges in simulating high-fidelity aerodynamic flows.
preprint: AIP/123-QED
I Introduction
Computational Fluid Dynamics (CFD) plays a pivotal role in solving and analyzing fluid flow problems, serving as an indispensable tool across diverse fields such as aerospace engineering, environmental science, industrial design, and biomedical research. Despite its broad applicability, a major challenge of CFD lies in its high computational cost. Techniques like Direct Numerical Simulation (DNS) and Large Eddy Simulation (LES) are computationally intensive, often making them impractical for many applications. While Reynolds-Averaged Navier-Stokes (RANS) models offer a more efficient alternative, they still fall short for real-time applications. These constraints limit CFD’s applicability in areas such as optimal flow control and airfoil optimization. To overcome these limitations, there is a critical need for the development of alternative methods and frameworks that improve computational efficiency without compromising simulation accuracy.
In recent years, the emergence of deep learning techniques has opened new avenues for efficient flow field predictions. A growing body of research has integrated deep learning methods with fluid flow simulation, yielding promising results. Notably, neural network architectures such as U-net Guo, Li, and Iorio (2016); Sekar et al. (2019); Ribeiro et al. (2020); Chen et al. (2021); Cai et al. (2022), generative adversarial networks (GANs) Goodfellow et al. (2020); Chen, Chiu, and Fuge (2020); Haizhou et al. (2022), and physics-informed neural networks (PINNs) Raissi, Perdikaris, and Karniadakis (2019); Cai et al. (2021); Chen, Liu, and Sun (2021); Zhang et al. (2023) have shown considerable potential in tasks such as flow field reconstruction. These approaches demonstrate the power of deep learning to capture the complex dynamics of fluid flow, signaling a shift in the future of CFD research. Guo et al.Guo, Li, and Iorio (2016) proposed a convolutional neural network (CNN)-based surrogate model for real-time prediction of steady laminar flow fields, achieving significantly faster velocity field estimation compared to traditional CFD solvers while maintaining low error rates. Raissi et al.Raissi, Perdikaris, and Karniadakis (2019) introduced physics-informed neural networks , a deep learning framework that integrates physical laws governed by nonlinear partial differential equations to solve forward and inverse problems, enabling data-efficient spatio-temporal function approximation and scientific discovery even in small data regimes.
The deep learning landscape is still rapidly evolving, with newer architectures—such as Transformers Vaswani (2017); Khan et al. (2022); Lin et al. (2022) and denoising diffusion probabilistic models (DDPMs) Ho, Jain, and Abbeel (2020); Yang and Sommer (2023, 2023)—along with scalable diffusion models integrated with Transformers (DiT) Peebles and Xie (2023), showing superior performance in tasks like image generation. This evolution raises an important question: Can the latest advancements in machine learning enable real-time, accurate predictions of flow fields, effectively addressing the challenges of computational efficiency and accuracy?
Although recent advancements using DDPMs have shown promising results in fluid dynamics prediction, these models have primarily been applied to relatively simple flow fields. For instance, methods like FluidDiff Yang and Sommer (2023) and the approach by Shu et al. Shu, Li, and Farimani (2023) have demonstrated success in prediction and reconstruction tasks, but they are generally limited to simple flow conditions. Furthermore, the time-consuming sampling process inherent in U-Net-based DDPMs makes them unsuitable for real-time predictions. These limitations highlight the urgent need for more efficient and practical methods capable of handling complex, real-time fluid field predictions.
Efficient and accurate prediction of high Reynolds number airflow around airfoils is crucial for advancements in aerodynamics research and the aviation industry, as it directly impacts airfoil optimization, optimal flow control, and the design of more efficient aerodynamic systems. The application of deep learning in aerodynamics has evolved progressively from foundational flow field prediction to efficient optimization design. Early research (Sekar, 2019) established a speed foundation for replacing traditional CFD through a hybrid CNN-MLP architecture (where CNNs extract geometric features and MLPs map flow parameters), enabling real-time flow field prediction, though limited to steady laminar flows. To address data dependency issues, Wu (2021) introduced a Generative Adversarial Network (daGAN), leveraging a two-phase training mechanism (cGAN pre-training followed by dual-generator fine-tuning) to overcome sparse data constraints and enhance generalization under small-sample conditions. Li (2020) developed Deep Learning-based Geometric Filtering (DLGF), utilizing autoencoder-based compression of the design space to reduce variable dimensionality from hundreds to tens, thereby circumventing the curse of dimensionality and accelerating optimization workflows. Within this framework, Wang (2022) integrated a CNN-based Reduced Order Model (ROM) with a fusion strategy of pooling dimensionality reduction and radial basis functions to handle nonlinear flow optimization under large disturbances. Concurrently, Li (2021) addressed low-Reynolds-number separation bubbles by integrating the XFOIL solver with a transition-sensitive turbulence model, achieving efficient airfoil design for UAVs through tailored modal decomposition.
While these works mentioned above collectively aim to replace high-cost simulation iterations via a paradigm integrating physical constraints with data-driven intelligence, they exhibit limitations: conventional generative models (e.g., GANs, DDPMs) inadequately capture complex separated flows in high-Reynolds-number turbulence; U-Net-based diffusion models suffer from inefficient iterative sampling that precludes real-time application; and purely data-driven approaches show weak extrapolation capability under extreme Reynolds numbers while lacking embedded physical constraints, leading to localized errors near critical airfoil surfaces. However, due to the complexity of the flow—characterized by turbulence and flow separation—achieving both efficiency and accuracy in these predictions remains a significant challenge.
This study introduces a novel methodology for aerodynamic flow-field simulation—Adaptive Diffusion Transformers for Airfoil Flow Simulation (AeroDiT)—which employs diffusion transformers to learn from high-fidelity Reynolds-Averaged Navier–Stokes (RANS) simulation data. By treating the complex flow-field distributions as high-dimensional probability spaces and leveraging a generative approach, AeroDiT aims to dramatically reduce computational costs while preserving predictive accuracy at high Reynolds numbers. Compared to conventional computational fluid dynamics (CFD) techniques, AeroDiT’s innovations extend both the theoretical and practical frontiers of aerodynamic simulation, yielding the following key contributions:
1. Pioneering Application of Diffusion Transformers in Aerodynamic Flow Prediction. This work represents the first utilization of Scalable Diffusion Models with Transformers (DiTs) as surrogate models for predicting aerodynamic flow fields. Traditional CFD approaches to high-Reynolds-number turbulence require fine spatial discretization and iterative solvers, resulting in prohibitive computational overhead. AeroDiT reframes the flow-field as a high-dimensional latent distribution and embeds multiscale turbulent structures into a diffusion-based generative framework. The transformer’s self-attention mechanism captures interactions across all spatial scales, enabling efficient reconstruction of the entire flow field from learned statistical representations. From a theoretical standpoint, this extends the application paradigm of generative models within physical sciences and demonstrates a data-driven route to approximately solve partial differential equations (PDEs), thereby circumventing the heavy numerical discretization typical of classical methods.
2. Experimentally Validated, High-Fidelity, Real-Time Prediction Capability. During validation, AeroDiT exhibits exceptional accuracy and real-time performance. Across a variety of angles of attack, freestream velocities, and representative airfoil geometries, the model consistently achieves relative L2 prediction errors on velocity fields, pressure distributions, and turbulence metrics (such as vorticity) on the order of only a few percent. Simultaneously, the generative inference process can reconstruct a complete flow field in a matter of seconds—many orders of magnitude faster than RANS-based CFD. Consequently, AeroDiT holds immense promise for applications requiring massive parametric sweeps, real-time control, or rapid design iteration. It can accelerate the design optimization cycle in aerospace engineering and furnish reliable surrogate models for data-driven, multidisciplinary design optimization.
3. Integration of Physical Priors to Ensure Consistency with Governing Equations. Beyond empirical accuracy and speed, AeroDiT’s core innovation lies in blending generative modeling with physics-informed constraints. By embedding RANS-derived physical priors during training, the network continually corrects deviations from conservation laws, preventing unbounded extrapolation errors in out-of-distribution regimes. Such a methodology paves the way for extending generative deep learning to more complex flow problems (e.g., multiphase flows, combustion, three-dimensional wing arrays), offering a general framework for achieving a balance between physical fidelity and statistical expressivity.
The remainder of the paper is organized as follows: Section II reviews related work our work has been built on, and provide an overview of the background information. Section III outlines the fundamental methodology used in this work. Section IV presents the experimental evaluation, detailing datasets, configurations, and results obtained through extensive testing. Section V discusses the implications of the results, highlighting the strengths and limitations of the proposed approach while suggesting future research directions.
II Related Work
II.1 Diffusion Models in Fluid Dynamics
Recent advances in deep learning-based fluid dynamics prediction have utilized generative models to address the challenges of simulating complex flow fields. Yang et al. Yang and Sommer (2023) introduce FluidDiff, a novel generative model based on Denoising Diffusion Probabilistic Models (DDPMs). FluidDiff employs a diffusion process to learn a high-dimensional representation of dynamic systems and uses Langevin sampling to predict flow states under specific initial conditions. Trained on finite, discrete fluid simulation data, FluidDiff achieves competitive performance, accurately predicting test data without the need for explicit encoding of the underlying physical system’s prior knowledge. Similarly, Shu et al. Shu, Li, and Farimani (2023) propose a DDPM-based model for flow field super-resolution, trained on high-fidelity data to reconstruct detailed flow fields from low-fidelity or sparsely sampled data. This model also incorporates physics-informed conditioning through partial differential equations (PDEs), improving its predictive accuracy. Hu et al. Hu, Lu, and Yang (2024) introduce the Geometry-to-Flow (G2F) diffusion model, which generates flow fields around obstacles by conditioning on obstacle geometry. Zhou et al. Zhou et al. (2024) present Text2PDE, a latent diffusion model framework for PDE simulation that uses autoencoders and text-based conditioning to generate full spatio-temporal solutions. Liu et al.Liu and Thuerey (2024) proposed an uncertainty-aware surrogate model for turbulence simulations using DDPMs, demonstrating superior accuracy and uncertainty quantification compared to Bayesian neural networks and heteroscedastic models. While these studies show promising results, many of them primarily focus on simpler flow fields, leaving room for further exploration in more complex and high-dimensional flow prediction scenarios. To address these limitations, there is increasing interest in developing more advanced architectures capable of modeling such complexities, such as the Diffusion Transformer, which is introduced in the next section.
II.2 Diffusion Transformer
In recent years, Transformers have achieved remarkable success across various domains, including natural language processing and computer vision Vaswani (2017). Meanwhile, diffusion models, as a class of generative models, have shown significant potential in image generation tasks. The Diffusion Transformer (DiT) combines the strengths of Transformers with a diffusion-based architecture, demonstrating outstanding performance in diverse applications Peebles and Xie (2023). Despite these advances, the application of DiT models to aerodynamic simulation remains unexplored. To the best of our knowledge, this study represents the first application of a DiT-based model trained on Reynolds-averaged Navier-Stokes (RANS) simulation data, with a particular focus on predicting flow around airfoils. In addition, in line with prior DiT-based work Ovadia et al. (2023), our study addresses the challenge of real-time inference. A detailed description of the methodology is provided in the following section.
III Methodology
Figure 1: The framework and learning process of the proposed AeroDiT (Normalized flow field of EPPLER 59 airfoil).
This study employs Diffusion Transformers to predict the airfoil flow field at high Reynolds numbers. The framework of Diffusion Transformers (DiTs) is illustrated in Figure 1. By leveraging the generative capabilities of diffusion models, DiTs function as a surrogate model that effectively learns the underlying statistical distribution of flow field data. The model is designed to predict the flow characteristics, such as velocity components and pressure fields, based on input features including the geometry of the airfoil and various flow conditions. The overall architecture consists of three primary components: the Condition Encoder, the Diffusion Transformer (DiT) block, and the Function Space Decoder. The Condition Encoder processes input data such as the airfoil geometry and flow velocities to generate an initial latent representation. This encoded information is then fed into the DiT block, where noise is added, and a series of linear transformations, reshaping, layer normalization, and attention mechanisms take place to effectively model the underlying data distribution. The DiT block utilizes multi-head self-attention, pointwise feed-forward networks, and residual connections to capture the complex relationships between the variables in the latent space. Finally, the Function Space Decoder decodes the learned representation into meaningful output data, such as the predicted pressure field and velocity components (both horizontal and parallel), which are visualized in the final output.
Once trained, DiTs can be used as a predictive tool, providing accurate approximations of the flow field for different aerodynamic configurations, without the need for traditional computational fluid dynamics (CFD) simulations. This offers a computationally efficient and scalable alternative for real-time flow field predictions based on airfoil geometry and specified flow conditions. The geometric shape of the wing is shown in Figure 2, with blue for the flow field region and red for the wing shape.
The training process consists of two key steps: forward diffusion and reverse denoising, as shown in Figure 3. In the forward diffusion step, Gaussian noise is progressively added to the input flow field, gradually transforming it into a noise distribution. The reverse denoising process then reconstructs the original flow field iteratively, utilizing Bayesian inference to model the transition probabilities. Together, these processes enable the model to effectively approximate the underlying distribution and generate accurate predictions. The following sections introduce the relevant notations and learning objectives, followed by a detailed description of the forward and reverse diffusion processes.
III.1 Notations
To ensure clarity, we define the key symbols used throughout this study:
  • $\bm{x}$: Flow field, including pressure $\mathbf{p}$ and velocity components $u_{x}$ (along $x$-axis) and $u_{y}$ (along $y$-axis).
  • $\bm{\Psi}$: Input conditions, including airfoil geometry and inflow conditions $u_{x}$, $u_{y}$.
  • $\mathbf{p}$: Pressure field. Here, we bold $\mathbf{p}$ to distinguish it from the probability distribution $p$.
  • $\mathbf{S}(\bm{\Psi})$: The AeroDiT surrogate model.
  • $\mathbf{R}(\bm{\Psi})$: The real physical system.
  • $\epsilon\sim\mathcal{N}(0,\mathbf{I})$: Gaussian noise added during forward diffusion.
  • $\alpha_{t},\bar{\alpha}_{t}$: Variance scaling factors in the diffusion process.
III.2 Learning Target
The primary objective is to develop a surrogate model, $\mathbf{S}$, that approximates the real physical system, $\mathbf{R}$, governed by the Navier-Stokes equations. As discussed before, solving these equations numerically for real flow fields $\mathbf{R}(\bm{\Psi})$ is computationally expensive. To address this, the learning target is defined as identifying the optimal parameters $\theta$ of the surrogate model $\mathbf{S}$ such that:
$\bm{x}=\mathbf{S}_{\theta}(\bm{\Psi})\approx\mathbf{R}(\bm{\Psi}),$ (1)
From the perspective of generative modelling, this can be achieved by learning the conditional distribution $p(\bm{x}|\bm{\Psi})$ from our dataset $\bm{D}$, which captures the statistical relationship between the input conditions $\bm{\Psi}$ and the resulting flow field $\bm{x}$. To approximate the real physical system, the model optimizes parameters $\theta$ by maximizing the likelihood of observed flowfield data:
$\max_{\theta}\mathbb{E}_{\bm{x},\bm{\Psi}}\big{[}\log p_{\theta}(\bm{x}|\bm{\Psi})\big{]}.$ (2)
Therefore we formulate our learning target as learning the conditional distribution $p(\bm{x}|\bm{\Psi})$ from our dataset $\bm{D}$. By learning this distribution, the model can generate accurate predictions efficiently, bypassing the computational complexity of solving the Navier-Stokes equations numerically. To approximate such distribution, our AeroDiT model goes through a forward diffusion process and a reverse denoising process, as discussed in the following sections.
Figure 2: Wing geometry. (The red part shows the wing and the blue part shows the flow field)
Figure 3: Illustration of the forward and backward diffusion process.
III.3 Forward Diffusion Process
In the forward diffusion process, we followed the canonical procedures described in literature Ho, Jain, and Abbeel (2020); Sohl-Dickstein et al. (2015). As shown in Figure 3, we gradually add noise to the data $\bm{x}_{0}$ (the true flow field) over several time steps given the input conditions $\bm{\Psi}$, transforming it into a noisy sample $\bm{x}_{T}$. The noise schedule $\beta_{t}$ is pre-defined and fixed, increasing with each timestep. At each timestep $t$, the conditional probability for adding noise is given by:
$q(\bm{x}_{t}|\bm{x}_{t-1},\Psi)$ $=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\mathbf{I})$ (3)
where $\bm{x}_{t}$ represents the flow field at timestep $t$, and $\beta_{t}$ is the pre-defined noise variance at that timestep. The noise is added progressively until, at the final timestep $T$, the data becomes almost entirely noise.
The main goal of this process is to transform the data into a noise distribution that can be efficiently learned by the reverse denoising process. Since the variable $\bm{\Psi}$ does not break the Markovian nature of the forward diffusion process, the forward diffusion process for $t\in[1,T]$ is formally described as:
$q(\bm{x}_{t}|\bm{x}_{0},\Psi)=\mathcal{N}(\bm{x}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (4)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ and $\bm{x}_{0}$ is the original clean data. The term $\bar{\alpha}_{t}$ is the cumulative product of the noise schedule, determining the level of corruption at each timestep. The noise becomes larger as $t$ increases, leading to a greater amount of information lost by timestep $T$. The denoising process will learn to remove the noise and hence learn the conditional distribution, as discussed in the next section.
III.4 Reverse Denoising Process and the Loss
In the reverse denoising process, we start with the noisy data $\bm{x}_{T}$ obtained from the forward diffusion process and iteratively remove the noise to recover the original data $\bm{x}_{0}$. The ultimate goal for the denoising process is to help achieve our learning target discussed in Section III.2.
By constructing the forward diffusion process, we can employ variational inference to show that minimizing the Kullback-Leibler (KL) divergence
$KL\big{(}q(\bm{x}_{t-1}\mid\bm{x}_{t},\bm{\Psi})\,\|\,p_{\theta}(\bm{x}_{t-1}\mid\bm{x}_{t},\bm{\Psi})\big{)}$
effectively maximizes the Evidence Lower Bound (ELBO) of
$\mathbb{E}_{\bm{x},\bm{\Psi}}\big{[}\log p_{\theta}(\bm{x}|\bm{\Psi})\big{]}.$
Here, the KL divergence measures how closely two probability distributions align. This insight allows us to reformulate the learning target: instead of directly maximizing $\mathbb{E}_{\bm{x},\bm{\Psi}}\big{[}\log p_{\theta}(\bm{x}|\bm{\Psi})\big{]}$ we focus on learning the optimal parameters $\theta$ so that $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{\Psi})$ accurately approximates $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{\Psi})$. In other words, by correctly learning the denoising process, we approximate the conditional distribution in a stepwise manner.
The reverse denoising process $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{\Psi})$ follows a Gaussian distribution, with its mean and variance derived asHo, Jain, and Abbeel (2020):
$\tilde{\mu}_{t}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}\right)$ (5)
$\tilde{\beta}_{t}$ $=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ (6)
We notice that $\tilde{\beta}_{t}$ depends only on the predetermined hyperparameters, and we assign such a constant value directly to the variance of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{\Psi})$. This leaves us only with the need to adjust $\theta$ so that the mean converges. By observation, one can realize that making $\mu_{\theta}$ converge to $\tilde{\mu}$ is equivalent to making the noise prediction converge to the noise added in the diffusion process. Therefore, from our initial learning target, we formulate our final learning loss:
$\mathcal{L}_{MSE}(\theta)=\mathbb{E}_{t,\bm{x}_{0},\bm{\Psi},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\bm{x}_{t},t,\bm{\Psi})\|^{2}\right]$ (7)
In the DiT framework, the noise term, $\epsilon_{\theta}(\bm{x}_{t},t,\bm{\Psi})$, is trained and predicted using a Transformer-based architecture, enabling the model to capture long-range dependencies and contextual information. This contrasts with DDPMs, which typically employ simpler architectures, such as MLPs or CNNs. By leveraging self-attention, DiT incorporates richer contextual information, improving its ability to capture dependencies and correlations within the data. This enhances the model’s ability to handle complex inputs, such as structured data or sequential dependencies (e.g., time series or flow fields). Rather than directly predicting the noise, DiT predicts a denoised version of the flow field $x_{0}$, utilizing attention mechanisms. Although the predicted noise $\epsilon_{\theta}$ is still required, it is inferred in a manner distinct from that in DDPM-based models.
By accurately predicting the noise, the model learns the stepwise transformations required to approximate the conditional distribution $p(\bm{x}\mid\bm{\Psi})$. Based on this theoretical foundation, the following section focuses on the practical implementation of these concepts.
III.5 Overall Training
As shown in Figure 1, the model is trained in a sequential manner, starting with the encoder, followed by the decoder, and finally the DiT network. Specifically, an asymmetric training approach is employed for the conditional encoder and the functional space decoder.
The conditional encoder aims to extract task-relevant features from the input conditional data. By incorporating specific conditional information, it generates conditionally enriched feature representations, which are then fed into the DiT network to guide the generation of conditional output features in the diffusion process.
The functional space decoder plays a pivotal role in mapping the features produced by the DiT network to the target task space, effectively achieving the specific task objectives for data generation. During feature decoding, it leverages task-specific properties to enhance the model’s generalization performance.
To train DiT module, we employ the MSE loss $\mathcal{L}_{MSE}(\theta)$ to quantify the discrepancy between the noise predicted by DiT module and the true noise. The DiT training procedures are summarized in Algorithm 1.
Input: Arguments args (batch size, epochs, etc.), Trained VAE encoder-decoder.
Output: Trained DiT model
1
2Initialize distributed training environment;
3
4Initialize DiT model, EMA model, optimizer, and trained VAE models;
5
[…]
7       for each batch (condition, physical field) $(x,y)$ in DataLoader do
8
9            Encode $x$ and $y$ into the latent space using VAE models;
10
11            Sample random time steps $t$;
12
13            Compute $\mathcal{L}_{MSE}$ loss;
14
15            Backpropagate and update DiT model parameters;
16
17       end for
18
19 end for
20
Algorithm 1 Training DiT Model
With the training procedures described above, the next section focuses on the experimental evaluation.
IV Experimental Evaluation
To evaluate the performance of AeroDiT, we conducted extensive experiments using real-world datasets. The following sections detail the experimental setup, present a discussion of the results, and delve into specific case studies for a more in-depth analysis. All the source codes to reproduce the results are available on https://github.com/AI4SciFoundation/DiT4Science.
IV.1 Experiment Settings
IV.1.1 Datasets and Data preparation
In this study, we utilize three datasets and a test set, as described in Thuerey et al. (2020). These datasets are generated using RANS simulations to compute the velocity and pressure distributions of incompressible flows around airfoils. The simulations cover a wide range of Reynolds numbers ($Re=[0.5,5]$ million) and angles of attack (±$\pm 22.5^{\circ}$), providing a diverse set of aerodynamic conditions. A total of 1505 unique airfoil shapes were sourced from the UIUC database and combined with randomly sampled freestream conditions to create the training inputs. The RANS solutions were computed using the Spalart-Allmaras (SA) turbulence model, a widely used one-equation model for industrial applications. The computations were carried out with OpenFOAM, an open-source CFD software.
The first dataset is a regular dataset (Regular), which contains 220,000 images with a data size of (6, 128, 128). The six channels in each image represent: the initial velocity along the $x$-axis (channel 1), the initial velocity along the y-axis (channel 2), the airfoil mask (channel 3), pressure (channel 4), velocity along the $x$-axis (channel 5), and velocity along the $y$-axis (channel 6). The primary task of our approach is to train the model to correctly predict the pressure, $x$-axis velocity, and $y$-axis velocity based on the first three channels. The second dataset is an augmented dataset (Shear), which also consists of 220,000 images. It is generated by shearing the wing shape along the center axis by ±15 degrees to expand the shape space seen by the network, providing a richer variety of inputs. The final dataset is a small subset of the regular dataset (Subset), containing 6,000 images. This subset is introduced to better assess the model’s generalization ability when trained with a limited amount of data.
The test set comprises 90 images, which do not appear in the training data and remain unseen during all stages prior to the testing phase. The Reynolds number distribution in the test set spans a wide range, from [$2.04\times 10^{6}$ to $2.00\times 10^{7}$], encompassing 30 different airfoil types. All the datasets used in this study are publicly available TUM-I Repository 1 (2018); TUM-I Repository 2 (2018).
IV.1.2 Configuration
We employ the DiT (Scalable Diffusion Models with Transformers Peebles and Xie (2023)) model as backbone for AeroDiT and adopt the same encoder and decoder models as in Ho, Jain, and Abbeel (2020). We use three configs: DiT-B, DiT-L and DiT-XL. Initially, we train two separate VAE-based encoder-decoder models for 100 epochs:
  • Condition encoder-decoder: Input is a tensor of shape (3, 128, 128), where the channels represent initial x-velocity, initial y-velocity, and the airfoil mask.
  • Function space encoder-decoder: Input is a tensor of shape (3, 128, 128), with channels representing pressure, x-velocity, and y-velocity.
After training, the conditional encoder is retained as the encoder, and the spatial decoder is retained as the decoder. The trained encoder and decoder are used to train the DiT model for 2,500,000 iterations. The experimental results are presented as the average of three random tests. During the testing phase, unless otherwise specified, the number of sampling steps is set to 100, and the model performance is evaluated using the L2 norm error between the predicted and ground truth.
Table 1: L2 Error of Airfoil Flow Simulations Across Multiple Datasets.
DiT-B/2 DiT-L/2 DiT-XL/2
Dataset p ux uy p ux uy p ux uy
Subset $0.1927^{\pm 0.23\%}$ $0.0262^{\pm 0.02\%}$ $0.0865^{\pm 0.03\%}$ $0.1989^{\pm 0.42\%}$ $0.0285^{\pm 0.14\%}$ $0.0909^{\pm 0.33\%}$ $0.1935^{\pm 0.13\%}$ $0.0271^{\pm 0.07\%}$ $0.0884^{\pm 0.05\%}$
Shear $0.2000^{\pm 0.39\%}$ $0.0470^{\pm 0.22\%}$ $0.1120^{\pm 0.30\%}$ $0.1964^{\pm 0.19\%}$ $0.0458^{\pm 0.13\%}$ $0.1116^{\pm 0.08\%}$ $0.1864^{\pm 0.76\%}$ $0.0415^{\pm 0.24\%}$ $0.0105^{\pm 0.37\%}$
Regular $0.1065^{\pm 0.58\%}$ $0.0251^{\pm 0.22\%}$ $0.0547^{\pm 0.22\%}$ $0.1041^{\pm 0.54\%}$ $0.0250^{\pm 0.23\%}$ $0.0537^{\pm 0.34\%}$ $0.1103^{\pm 0.26\%}$ $0.0278^{\pm 0.22\%}$ $0.0578^{\pm 0.31\%}$
IV.2 Results and Discussion
Table 1 summarizes the prediction results of the model using different DiT backbones. In experiments conducted on the annotated dataset, the best predicted $L_{2}$ errors were consistently below 0.105, 0.025, and 0.055, respectively, across all test cases. The “$\pm$” values indicate the relative error percentage with respect to the mean $L_{2}$ error, providing a measure of variability across different test samples.
We provide more specific details on the model training pro- cess. Figure 4(a) illustrates the distribution of Reynolds numbers in the test set, which are uniformly distributed over a wide range from $2\times 10^{6}$ to $2\times 10^{7}$. Figure 4(b) shows the variation of the experimental loss with the number of training iterations for the backbone network, revealing that the loss stabilizes after approximately $5\times 10^{5}$ iterations. Finally, Figure 4(c) demonstrates that the model’s performance on the test domain becomes stable after more than $2\times 10^{5}$ training iterations.
Figure 4(d) evaluates the model’s performance with varying sampling steps during testing. Notably, we observe that the prediction accuracy stabilizes after more than three sampling steps. At this point, the time required to predict a single instance is reduced to a matter of seconds, allowing our method to perform near-real-time predictions on the test set.
Figure 4: Further studies on AeroDiT training and inference process.
Training resource consumption. All experiments were conducted on an NVIDIA RTX 6000 GPU. Table 2 compares different model types of DiT (DiT-XL/2, DiT-L/2, DiT-B/2) in terms of GPU memory usage, total number of parameters and training speed (training steps per second). As the model size decreases (from DiT-XL/2 to DiT-B/2), the GPU memory usage significantly drops, while the training speed increases.
Table 2: Training resource consumption.
Model Type Parameters GPU Memory Usage Training Steps/Sec
DiT-XL/2 673,774,880 21,265 MiB 7.79
DiT-L/2 456,898,592 14,037 MiB 10.89
DiT-B/2 129,609,248 5,535 MiB 23.67
IV.3 Case Study
To evaluate the performance of AeroDiTs, we evaluated their performance by comparing predictions against CFD benchmarks across various airfoil geometries and Reynolds numbers, as shown in Table 3. Note that the selected airfoils and Reynolds numbers belong to the target domain, distinct from the source domain used for training. This ensures that the evaluation accurately reflects the model’s ability to generalize to unseen conditions. The chosen test cases are designed to assess the model’s generalization across a variety of airfoil shapes and a wide range of Reynolds numbers (2.04–20.00 million), which are representative of the complex turbulent flows encountered in real-world applications.
Table 3: Configuration of the airfoil and Reynolds number in the test case.
Test Case Airfoil Reynolds Number
1 Drela AG09 $4.04\times 10^{6}$
2 AH 63-K-127/24 $9.11\times 10^{6}$
3 EPPLER 59 $1.21\times 10^{7}$
Figure 5: Normalized flow field of Drela AG09 airfoil at $Re=4.04\times 10^{6}$.
Figure 6: Comparison of surface pressure and velocities of Drela AG09 airfoil at $Re=4.04\times 10^{6}$.
Figure 5 shows the normalized pressure and velocity field obtained by AeroDiTs and the CFD of the Drela AG09 airfoil at $Re=4.04\times 10^{6}$. Qualitatively, the predictions of AeroDiTs are in good agreement with CFD results. For the pressure field, the maximum absolute error of the prediction result is around 0.09, while for the velocity field in both directions, the corresponding prediction error does not exceed 0.010 and 0.041. It can be found that the positions with the largest prediction errors all occur on the airfoil surface. Therefore, in order to quantitatively assess the prediction error, we give the flow field results along the airfoil surface, as shown in Figure 6. Overall, the longitudinal velocity along the airfoil surface is best predicted, with some differences in surface pressure and transverse velocity, but the overall trends of the curves are consistent.
The prediction errors on the airfoil surface are primarily concentrated in the pressure and transverse velocity fields, which are likely related to flow separation and boundary layer characteristics. In high Reynolds number flows, the turbulent structures and separation phenomena in the boundary layer significantly influence the flow field. AeroDiT may struggle with these complex flow features. From a physical perspective, these errors reflect the model’s limitations in capturing the fine details of the flow.
Moreover, AeroDiT achieves rapid predictions through its diffusion transformer architecture, significantly enhancing computational efficiency. However, this efficiency gain comes at the cost of some accuracy loss. While the model excels in predicting longitudinal velocity, errors persist in the pressure and transverse velocity predictions. This underscores the need to balance computational efficiency and prediction accuracy when designing deep learning models. For practical engineering applications, this trade-off should be adjusted according to specific needs. For instance, efficiency can be prioritized in the preliminary design phase, while higher accuracy is required in the optimization phase.
V Discussion
This study demonstrates the significant potential of the Diffusion Transformer in improving airflow simulations for high Reynolds number airfoil flows. We evaluated AeroDiT against traditional CFD results using three distinct airfoil types and Reynolds numbers ranging from 2 million to 20 million, which were not included in the training dataset.
For the Drela AG09 airfoil at a Reynolds number of $4.04\times 10^{6}$, AeroDiT showed strong agreement with CFD, with a maximum absolute error of approximately 0.02 in the pressure field and no more than 0.06 in the velocity field. In a more challenging test case with the AH 63-K-127/24 airfoil at a Reynolds number of $9.11\times 10^{6}$, AeroDiT performed even better, with nearly all errors in the pressure and velocity fields remaining below 0.02. However, for the EPPLER 59 airfoil at a Reynolds number of $1.21\times 10^{7}$, performance was less satisfactory, likely due to the scarcity of high Reynolds number cases in the training dataset and the significant changes in flow field characteristics at this scale.
These findings highlight the effectiveness of DiT in generalizing across different airfoil shapes and Reynolds numbers, particularly in complex turbulence scenarios common in engineering applications. The study underscores the promise of integrating deep learning with traditional CFD methods, offering a novel approach to address long-standing challenges in accurately simulating airfoil flows.
Future directions could involve expanding the training dataset to include a broader range of high Reynolds number scenarios and incorporating physical laws into the learning process to enhance the model’s robustness and accuracy in more complex flow fields. Further research will also focus on developing domain generalization techniques Ma et al. (2023); Wang et al. (2022) based on varying Reynolds numbers. By incorporating precise Reynolds numbers as conditional inputs, we aim to further improve the predictive performance of the framework.
VI Conclusion
This study introduces AeroDiT, a novel framework that applies Diffusion Transformers to make real-time, accurate flow field predictions. By leveraging DiTs’ ability to learn conditional distributions $p(\bm{x}|\bm{\Psi})$, AeroDiT achieves reliable predictions of pressure and velocity fields, with average L2 errors of 0.1 for pressure ($p$) and 0.03 and 0.05 for velocity components ($u_{x}$, $u_{y}$), respectively. The results demonstrate the model’s ability to generalize across a broad range of Reynolds numbers ($2.04\times 10^{6}$ to $2.00\times 10^{7}$) and diverse airfoil geometries. In test cases involving the Drela AG09 and AH 63-K-127/24 airfoils, AeroDiT shows excellent agreement with CFD benchmarks, with prediction errors for pressure and velocity fields consistently below 0.02, even under challenging conditions.
Notably, AeroDiT exhibits strong performance in near-real-time scenarios, with the sampling steps converging to stable predictions within seconds. The transformer-based architecture significantly reduces computational costs, making high Reynolds number aerodynamic simulations more feasible for real-time applications. However, in cases involving extreme Reynolds numbers ($>1.2\times 10^{7}$), the model’s performance shows some limitations, likely due to the scarcity of corresponding training data and the drastic changes in flow field characteristics at these scales.
These findings highlight AeroDiT’s potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Looking ahead, the continued development and refinement of AeroDiT hold the promise of even greater advancements. By enriching the training dataset with higher Reynolds number scenarios and integrating physical constraints, we can further enhance the model’s performance in complex turbulence regimes. This not only paves the way for more accurate simulations but also opens new avenues for exploring the intricate dynamics of airflow around airfoils.
As we stand on the cusp of a new era in computational fluid dynamics, AeroDiT stands as a testament to the transformative power of machine learning. Its potential applications in various fields, from aerospace engineering to renewable energy, are vast and far-reaching. By bridging the gap between computational efficiency and predictive accuracy, AeroDiT heralds a future where real-time aerodynamic simulations are not just a possibility but a reality.
Appendix A ref figures
Figure 7: Distribution of pressure and velocity fields around the airfoil from low to high Reynolds numbers for the same airfoil : $Re_{1}=9.11\times 10^{6}$, $Re_{2}=1.06\times 10^{7}$, $Re_{3}=2.00\times 10^{7}$
Figure 8: Distribution of pressure and velocity fields around different airfoils at the same high Reynolds number : $Re=1.21\times 10^{7}$.